MICROWAVE SPECTROSCOPIC INVESTIGATIONS OF THE C–H⋯ π CONTAINING COMPLEXES CH_{2}F_{2} ⋯PROPYNE AND CH_{2}ClF⋯PROPYNE

\underline{REBECCA A. PEEBLES}, SEAN A. PEEBLES, CORI L. CHRISTENHOLZ, ANTHONY A. ERNST, and YASSER J. DHAHIR, Department of Chemistry, Eastern Illinois University, 600 Lincoln Ave., Charleston, IL 61920.

The spectra of the CH_{2}F_{2} ⋯propyne and CH_{2}ClF⋯propyne complexes have been studied by chirped-pulse and resonant cavity Fourier-transform microwave spectroscopy and by ab initio calculations at the MP2/6-311++G(2d,2p) level. Both complexes contain C–H⋯ π contacts, with the halogen atoms angled towards the methyl group end of the propyne. While CH_{2}F_{2} ⋯propyne has C_{s} symmetry, CH_{2}ClF⋯propyne has C_{1} symmetry, with the fluorine and chlorine atoms straddling the propyne.

Investigation of four single ^{13}C and the DC≡CCH_{3} isotopologues in CH_{2}F_{2} ⋯propyne has allowed a detailed structural determination, while only the ^{35}Cl and ^{37}Cl isotopologues have so far been assigned for CH_{2}ClF⋯propyne. Experimental data will be compared with ab initio results and with the analogous acetylene complexes, both of which have C_{s} symmetry structures, with double C–H⋯ π interactions.